1.     The time for a major exam to be completed is normally distributed with an average of 55 minutes and a standard deviation of 9 minutes. If 92% of the students completed the exam, when should the test be terminated?

 



1
Expert's answer
2021-11-08T19:47:50-0500

"P(Z<z)=0.92\to z=1.405."

"\frac{x-55}{9}=1.405\to x=1.405*9+55\approx 68."

The test should be terminated after 68 minutes.